How many hydrogen atoms are in 2.5 liters of ammonia gas at 760 torr and 27 degrees celcius?
1
Expert's answer
2017-04-07T10:16:07-0400
pV = nRT
n = PV/RT
760 torr = 760 mm = 1.0 atm
27°C = 27 + 273 K = 300 K
R = 0.0821 L atm / K mol
n = (1.0 atm x 2.5 L) / (0.0821 L atm / K mol x 300 K) = 0.1015 mol NH3
Nmolecules (NH3) = NA x n = 6.02 x 1023 x 0.1015 = 6.11 x 1022
N (H) = 6.11 x 1022 x 3 = 18.33 x 1022


Answer: N (H) = 18.33 x 1022